15,12,11,12,14,3,13,15. What is the mean and the median and round the answer to the naerest tenth.
Answer #1

(a)

The mean is given by:

\frac{\sum_{i=1}^{8}x^{i}}{8}=\frac{95}{8}=11.875

So,

Answer is:

11.875

(b)

Arranging numbers in ascending order, we get:

3,11,12,12,13,14,15,15

Median = (8+1)/2th item = Average of 4th & 5th item = (12 + 13)/2 = 12.5

So,

Answer is:

12.5